Charles is 23 years older than his sister carol. In 16 years he will be twice as old as carol. How old is each of them now?
Answer by ikleyn(52778)

.
Let H stands for Charles, C stands for Carol.

From the condition, you have these two equations:

H = C + 23,              (1)    ("Charles is 23 years older than his sister carol")

H + 16 =  2*(C+16).      (2)    ("In 16 years he will be twice as old as carol")


Substitute (1) into eq(2), replacing H. You will get a single equation for C:

(C+23) + 16 = 2*(C+16)  ====>

C + 39 = 2C + 32  ====>  39 - 32 = C  ====>  C = 7.


Answer.  Carol is 7 years old now.  Charles is 7+23 = 30 years old.


------------------
There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - Age problems for mental solution
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.